# KompeteAI: Accelerated Autonomous Multi-Agent System for End-to-End Pipeline Generation for Machine Learning Problems

### v1

@@ 3.3 Tree-Guided Exploration
    A set of FE node pairs $(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$ is sampled, excluding those present in the long-term memory buffer $\mathcal{M}_{\mathrm{long}}$. Each valid pair is merged into a new node
    $v_{ij}^{\mathrm{FE}}=\texttt{MergeFE}(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$
    which recombines structural and statistical traits of its parents.
@@
    For each merged node $v_{ij}^{\mathrm{FE}}$, a set of child MT nodes is generated from a conditional distribution. Then, for each parent FE node $v_{i}^{\mathrm{FE}}$ and $v_{j}^{\mathrm{FE}}$, the agent selects additional MT nodes from their respective subtrees. These are sampled stochastically with probabilities proportional to their scores:
    $u_{k}^{(i)}\sim\texttt{SampleTop}(v_{i}^{\mathrm{FE}})$
    The final child set of $v_{ij}^{\mathrm{FE}}$ combines the freshly generated MT nodes and the resampled top performers from its parents.
@@
    Additionally, a subset $\mathcal{V}_{\mathrm{FE}}^{\mathrm{merge}}\subseteq V_{t}^{\mathrm{FE}}$ is selected, and within each selected FE node, MT child pairs are merged using
    $u_{ij}^{\mathrm{MT}}=\texttt{MergeMT}(u_{i},u_{j})$
    to form stronger model configurations.
@@
Merging procedure detailed in Appendix A2.
3.4 Scoring Model
@@
The core principle behind the scoring model is to predict the final performance of a candidate model based on how similar models have performed on the same dataset. To achieve this, we generate detailed descriptions for several anchor ideas as well as for the candidate idea to be scored. This step is essential, as even minor code differences — such as the number of training epochs — can lead to significant variations in model performance. These descriptions are then inserted into a prompt to LLM with dataset description and current model training specification.
@@ 4.1 Motivation
While MLE-Bench is a significant step forward, it has two key limitations that hinder its practical utility and reproducibility. First, it is prohibitively large: even the ”Lite” version with just 22 competitions takes up 158GB and requires substantial compute to process. Second, MLE-Bench constructs test sets by partitioning the original training data and compares results to private leaderboard positions. This introduces significant evaluation bias, as it does not reflect actual test set performance. We show that scores on MLE-Bench’s constructed test sets can diverge significantly from real leaderboard rankings, leading to misleading conclusions and undermining its reliability as a proxy for real-world outcomes.
@@ 4.2 Methodology
  • Selection of Established Competitions. We curate 15 competitions from the ’lite’ MLEBench collection that still accept late submissions on Kaggle and whose individual dataset sizes do not exceed 1 GB. The aggregate volume of this subset is 5.3 GB, providing a stable foundation for baseline comparisons.
  • Incorporation of New Competitions. To better reflect the evolving landscape of AutoML tasks, we include 11 additional competitions from 2024 and 2025 years. These datasets, totaling 4.9 GB, were selected primarily to ensure fair comparison with human performance. This is because, given the recency of these competitions, both human participants and current models have access to almost the same tools and libraries, minimizing discrepancies due to technological advancements.
@@ 5.1 Experiment Setup
Baselines
To provide a comprehensive evaluation on MLE-Bench, we compared our system against the top-ranked methods on the leaderboard: AIDE, RD-agent, and Ml-Master. Due to the high cost of running all baselines, we report their leaderboard metrics as provided directly by MLE-Bench. For evaluation on Kompete-bench we re-ran only AIDE and RD-agent, as both are open-source, while the implementation of Ml-Master was not publicly available at the time of our study and was therefore excluded. All methods were tested using the same LLM backend — gemini-2.5-flash (Comanici et al. 2025), which also powers our system. Additionally, to fairly assess the alignment between medal distributions on MLE-Bench and real Kaggle leaderboards, we also evaluated all methods using o1-preview (Jaech et al. 2024), the model on which most MLE-Bench submissions achieved their highest scores.
Environment
To evaluate our system on MLE-Bench, we set a 6-hour runtime limit using 12 vCPUs, 64 GB of RAM, and a single NVIDIA A100 GPU (40 GB). The same hardware setup was used for all systems on Kompete-bench, where we applied a stricter 6-hour time limit. Each configuration was executed three times, and results were averaged to reduce variance. To ensure fairness and prevent systems from exploiting test-time information, we additionally RAG components from accessing any information about the competition date on which the agent was evaluated.
Benchmarks
We evaluate KompeteAI on the Lite subset of MLE-Bench, a computationally feasible proxy for the full benchmark. Although not identical, MLE-Bench reports high alignment in system rankings between Lite and full versions, making Lite a practical basis for comparison. For medal-based evaluation against real Kaggle leaderboards, we include only Lite competitions that still accept submissions. Additionally, we assess performance on our custom benchmark — Kompete-bench, structured following the same principles described in our methodology section, by separating the data into two categories: MLE-subset Lite and recent and a collection of newly curated competitions.
Evaluation Metrics
On MLE-Bench, we adopt the official leaderboard metric — the percentage of submissions that receive a medal — which aligns with standard Kaggle competition criteria. For Kompete-bench, we additionally report the percent humans beaten metric, which measures the percentage of human participants outperformed by the agent on the corresponding Kaggle leaderboard. This metric offers a more fine-grained and interpretable evaluation signal compared to medal thresholds, allowing us to distinguish between systems that may not earn medals but differ meaningfully in their relative competitiveness against human participants. To assess pipeline efficiency, we report the number of complete pipeline passes, where a single pass corresponds to one adding and one merging iteration performed by the AutoML system.
@@ 5.2 Results
Evaluation on MLE-Bench
We begin by evaluating our system on the widely adopted MLE-Bench benchmark (Lite subset). As presented in Table 2, KompeteAI attains the highest overall performance, outperforming existing state-of-the-art systems by an average margin of 3 percentage points.
@@
Table 2: Comparison of agents on MLE-Bench ’Lite’ subset. For AIDE and RD-agent, we use results reported by MLE-Bench. KompeteAI was run 3 times with different seeds; results are reported as mean ± SEM. Each run was given 6 hours.
Kaggle-Based Validation of Robustness
To validate robustness beyond proxy benchmarks, we tested all agents on the Lite subset using official Kaggle submissions (Table 3). The analysis reveals a substantial gap between MLE-Bench evaluations and actual leaderboard outcomes. In particular, both AIDE and RD-agent exhibit inflated medal rates on MLE-Bench compared to their real Kaggle performance. This discrepancy highlights the limitations of relying solely on MLE-Bench for comparison, especially in competitive human-level settings.
Name MLE-Bench LB Real LB Contemporary part
@@
Table 3: Medal rate gap between MLE-Bench leaderboard and leaderboard of real Kaggle competition. For this experiment we ran each system once and used the rules from Kaggle to determine the medal.
Evaluation on Kompete-Bench
To address these limitations, Kompete-Bench evaluates performance using real Kaggle leaderboards, covering recent competitions with high prize pools and strong engagement. On this Contemporary subset, traditional medal-based evaluation fails, as all agents achieve zero medal rates. We therefore propose a finer-grained metric: percent humans beaten. As shown in Figure 2, KompeteAI demonstrates state-of-the-art results, surpassing human performance in 11.2% of cases — significantly ahead of both RD-agent and AIDE. Nevertheless, the agent remains far behind top leaderboard teams, reflecting how contemporary competitions demand capabilities beyond local modeling — including large-scale feature engineering, synthetic data generation, and use of external datasets. Human participants leverage broader tools and resources, underscoring the competitiveness of these challenges and the remaining headroom for automated agents.
Impact of the Acceleration Methods
Finally, we assess the contribution of our acceleration framework. Table 4 reports the number of iterations within a fixed budget. By integrating predictive scoring and an accelerated debugging loop, our pipeline executes over 6.9× more iterations than a baseline without accelerations, enabling stronger submissions under identical constraints. Moreover, as shown in Figure 2, KompeteAI without accelerations exhibits a marked performance drop, underscoring its ability to refine solutions during test-time optimization. While we restrict acceleration measurements to KompeteAI due to tight pipeline integration, the proposed paradigms are general and may be adapted to other architectures with minimal conceptual changes.
Figure 2: Comparison of our pipeline with AIDE and RD-agent on Contemporary and MLE-Bench parts of Kompete-bench. All systems use gemini-2.5-flash as the underlying LLM. Each was run 3 times with different seeds, results are averaged, and each run was limited to 6 hours.
@@ 6 Conclusions
We introduced KompeteAI, a new multi-agent AutoML framework, and demonstrated its strong performance across diverse and challenging tasks. Our work opens several promising directions for future research.
@@
Acknowledgements
The study research has been supported by the Ministry of Economic Development of the Russian Federation (agreement No. 139-10-2025-034 dd. 19.06.2025, IGK  [^00] C313925P4D0002).
Appendix A Algorithms
A.1 Adding
@@ A.2 Merging
Appendix B Kompete-bench
Figure A3: End dates for competitions presented in Kompete-Bench
@@ Appendix B Kompete-bench
Appendix C Setup details
@@
KompeteAI
The list of hyperparameters used when running KompeteAI on benchmarks is given in the Table 6.The time_run_minutes parameter sets the maximum runtime for the entire multi-agent system in minutes, after which the process will terminate. The runtime_error_time defines the time limit (in minutes) after which a generated code is will be stopped. The subset_size_in_percent specifies the percentage of the dataset to be used for quick validation. The validator_size_threshold sets a threshold for the dataset size; if the data exceeds this value, a subset is used for training and validation. The number_of_ideas_eda determines how many exploratory data analysis (EDA) ideas generates per iteration. Similarly, number_of_ideas_data and number_of_ideas_modelling control the number of ideas related to feature engineering and model training. The max_add_idea parameter limits how many new ideas can be added to the idea pool in a single adding iteration. The number_of_selected_node specifies how many nodes are selected for adding expansion at each step. The number_of_iterations_parents sets how many iterations parent nodes participate in generating new ideas, while number_of_selected_node_merging determines how many nodes are chosen for merging at each iteration. The number_of_ideas_min and number_of_ideas_max define the minimum and maximum number of ideas, which are used as achor ideas for scoring model. The retrieve_n_papers and retrieve_n_competitions parameters control how many papers retrieves from arXiv papers and Kaggle solutions. The number_rag_ideas sets how many ideas are generated using RAG. The memory_size parameter determines how many recent ideas, solutions, or states each agent remembers for learning and decision-making. Alternatively, if memory_size is set to nearest_nodes, the agent’s memory consists of the most similar nodes or ideas, rather than a fixed number, allowing for more contextually relevant recall.
@@
RD-Agent
The list of hyperparameters used when running RD-agent on benchmarks is given in the Table 7. The debug_timeout parameter sets the maximum time, in seconds, that is allowed to debug one generated code. The full_timeout parameter defines the overall time limit for system. The if_action_choosing_based_on_UCB flag determines whether the agents select their actions using the Upper Confidence Bound (UCB) strategy. The enable_knowledge_base flag indicates whether a shared knowledge base is enabled for the agents. The loop_n parameter specifies the number of main iterations the system will execute, set to 2000 to avoid early stopping not by the time limit. Finally, the max_trace_num parameter limits the number of traces that can be created during system execution.
@@
AIDE
The list of hyperparameters used when running RD-agent on benchmarks is given in the Table 8. The steps parameter defines the maximum number of steps the entire system can perform during run. The max_debug_depth, specifies the maximum depth for recursive code debugging. The debug_prob parameter suggests that debugging is enabled for every generated code, meaning that all relevant information will be recorded without any sampling. Finally, the time_limit parameter represents the maximum allowed wall-clock time (in seconds) for experiment.
@@
Appendix D Metrics
Competition RD-agent AIDE KompeteAI aerial-cactus-identification 58 5 79 denoising-dirty-documents 3 None 12 dog-breed-identification 41 62 58 dogs-vs-cats-redux-kernels-edition 86 6 91 jigsaw-toxic-comment-classification-challenge 42 26 88 leaf-classification 38 24 46 mlsp-2013-birds None 0 0 nomad2018-predict-transparent-conductors 78 33 22 plant-pathology-2020-fgvc7 12 62 68 random-acts-of-pizza 58 47 67 spooky-author-identification 71 47 59 tabular-playground-series-dec-2021 17 22 25 tabular-playground-series-may-2022 16 32 36 text-normalization-challenge-english-language 3 9 27 text-normalization-challenge-russian-language 17 14 18

### v2

@@ 3.3.2 Merging
    A set of FE node pairs $(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$ is sampled, excluding those present in the long-term memory buffer $\mathcal{M}_{\mathrm{long}}$. Each valid pair is merged into a new node $v_{ij}^{\mathrm{FE}}=\texttt{MergeFE}(v_{i}^{\mathrm{FE}},v_{j}^{\mathrm{FE}})$ which recombines structural and statistical traits of its parents.
@@
    For each merged node $v_{ij}^{\mathrm{FE}}$, a set of child MT nodes is generated from a conditional distribution. Then, for each parent FE node $v_{i}^{\mathrm{FE}}$ and $v_{j}^{\mathrm{FE}}$, the agent selects additional MT nodes from their respective subtrees. These are sampled stochastically with probabilities proportional to their scores: $u_{k}^{(i)}\sim\texttt{SampleTop}(v_{i}^{\mathrm{FE}})$ The final child set of $v_{ij}^{\mathrm{FE}}$ combines the freshly generated MT nodes and the resampled top performers from its parents.
@@
    Additionally, a subset $\mathcal{V}_{\mathrm{FE}}^{\mathrm{merge}}\subseteq V_{t}^{\mathrm{FE}}$ is selected, and within each selected FE node, MT child pairs are merged using $u_{ij}^{\mathrm{MT}}=\texttt{MergeMT}(u_{i},u_{j})$ to form stronger model configurations.
@@
Merging procedure detailed in Appendix 2.
3.3.3 Scoring Model
@@
The core principle behind the scoring model is to predict the final performance of a candidate model based on how similar models have performed on the same dataset. To achieve this, we generate detailed descriptions for several anchor ideas as well as for the candidate idea to be scored. This step is essential, as even minor code differences — such as the number of training epochs — can lead to significant variations in model performance. These descriptions are then inserted into a prompt to LLM with the dataset description and current model training specification.
@@ 4.1 Motivation
While MLE-Bench is a significant step forward, it has two key limitations that hinder its practical utility and reproducibility. First, it is prohibitively large: even the ”Lite” version with just 22 competitions takes up 158 GB and requires substantial compute to process. Second, MLE-Bench constructs test sets by partitioning the original training data and compares results to private leaderboard positions. As shown in table 3, this introduces significant evaluation bias, as it does not reflect actual test set performance. We show that scores on MLE-Bench’s constructed test sets can diverge significantly from real leaderboard rankings, leading to misleading conclusions and undermining its reliability as a proxy for real-world outcomes.
@@ 4.2 Methodology
  • Selection of Established Competitions. We curate 15 competitions from the ’lite’ MLE-Bench collection that still accept late submissions on Kaggle and whose individual dataset sizes do not exceed 1 GB. The aggregate volume of this subset is 5.3 GB, providing a stable foundation for baseline comparisons.
  • Incorporation of New Competitions. To better reflect the evolving landscape of AutoML tasks, we include 11 additional competitions from the years 2024 and 2025. These datasets, totaling 4.9 GB, were selected primarily to ensure fair comparison with human performance. This is because, given the recency of these competitions, both human participants and current models have access to almost the same tools and libraries, minimizing discrepancies due to technological advancements.
@@ 5.1 Experiment Setup
Baselines. To provide a comprehensive evaluation on MLE-Bench, we compared our system with the top open-source methods on the leaderboard: AIDE, RD-agent, ML-Master, and MLE-STAR. Due to the high cost of rerunning all baselines, we report their leaderboard metrics as provided by MLE-Bench. For Kompete-Bench, all methods were tested with the same LLM backend — gemini-2.5-flash Comanici et al. (2025), which also powers our system, except ML-Master, which was run on deepseek-r1 Guo et al. (2025) due to architectural constraints. To ensure a fair comparison between MLE-Bench and Kompete-Bench and eliminate potential bias from using different LLM backends, we further evaluated AIDE and RD-Agent with o1-preview Jaech et al. (2024) — the model on which most MLE-Bench submissions achieved their highest scores.
Environment. To evaluate our system on MLE-Bench, we set a 6-hour runtime limit using 12 vCPUs, 64 GB of RAM, and a single NVIDIA A100 GPU (40 GB). The same hardware setup was used for all systems on Kompete-bench, where we applied a stricter 6-hour time limit. Each configuration was executed three times, and results were averaged to reduce variance. To ensure fairness and prevent systems from exploiting test-time information, we restrict RAG components in each run from accessing any information about the competition solutions, including those released after its completion.
Benchmarks. We evaluate KompeteAI on the Lite subset of MLE-Bench, a computationally feasible proxy for the full benchmark. Although not identical, MLE-Bench reports high alignment in system rankings between the Lite and full versions, making Lite a practical basis for comparison. For medal-based evaluation against real Kaggle leaderboards, we include only Lite competitions that still accept submissions. Additionally, we assess performance on our custom benchmark — Kompete-bench, structured following the same principles described in our methodology section, by separating the data into two categories: MLE-subset Lite and recent and a collection of newly curated competitions.
Evaluation Metrics. On MLE-Bench, we adopt the official leaderboard metric — the percentage of submissions that receive a medal — which aligns with standard Kaggle competition criteria. For Kompete-bench, we additionally report the percent humans beaten metric, which measures the percentage of human participants outperformed by the agent on the corresponding Kaggle leaderboard. This metric offers a more fine-grained and interpretable evaluation signal compared to medal thresholds, allowing us to distinguish between systems that may not earn medals but differ meaningfully in their relative competitiveness against human participants. To assess pipeline efficiency, we report the number of complete pipeline passes, where a single pass corresponds to one adding and one merging iteration performed by the AutoML system.
@@ 5.2 Results
Evaluation on MLE-Bench. We begin by evaluating our system on the widely adopted MLE-Bench benchmark (Lite subset). As presented in Table 2, KompeteAI attains the highest overall performance, outperforming existing state-of-the-art systems by an average margin of 3 percentage points.
@@
Table 2: Comparison of agents on MLE-Bench ’Lite’ subset. For AIDE, RD-agent, MLE-STAR and ML-Master, we use results reported by MLE-Bench. KompeteAI was run 3 times with different seeds; results are reported as mean ± SEM. Each run was given 6 hours.
Kaggle-Based Validation of Robustness. To validate robustness beyond proxy benchmarks, we tested all agents on the Lite subset using official Kaggle submissions (Table 3). The analysis reveals a substantial gap between MLE-Bench evaluations and actual leaderboard outcomes. In particular, all comparable systems exhibit inflated medal rates on MLE-Bench compared to their real Kaggle performance. This discrepancy highlights the limitations of relying solely on MLE-Bench for comparison, especially in competitive human-level settings.
Name MLE-Bench LB, % Real LB, % LBs medal intersection, %
@@
ML-Master (deepseek-r1) 27 13 7
MLE-Star (gemini-2.5-flash) 20 13 7
Table 3: Medal rate gap between the MLE-Bench leaderboard and the leaderboard of a real Kaggle competition. The ”LBs medal intersection” column indicates the percentage of competitions where a medal was won on both the real LB and the MLE-Bench LB. For this experiment we ran each system once and used the rules from Kaggle to determine the medal.
Evaluation on Kompete-Bench. To address these limitations, Kompete-Bench evaluates performance using real Kaggle leaderboards, covering recent competitions with high prize pools and strong engagement. On this Contemporary subset, traditional medal-based evaluation fails, as all agents achieve zero medal rates. We therefore propose a finer-grained metric: percent humans beaten. As shown in Figure 2, KompeteAI demonstrates state-of-the-art results, surpassing human performance in 14% of cases — significantly ahead of all comparable systems. Nevertheless, the agent remains far behind top leaderboard teams, reflecting how contemporary competitions demand capabilities beyond local modeling — including large-scale feature engineering, synthetic data generation, and use of external datasets. Human participants leverage broader tools and resources, underscoring the competitiveness of these challenges and the remaining headroom for automated agents.
Impact of the Acceleration Methods. Finally, we assess the contribution of our acceleration framework. Table 4 reports the number of iterations within a fixed budget. By integrating predictive scoring and an accelerated debugging loop, our pipeline executes over 6.9× more iterations than a baseline without accelerations, enabling stronger submissions under identical constraints. Moreover, as shown in Figure 2, KompeteAI without accelerations exhibits a marked performance drop, underscoring its ability to refine solutions during test-time optimization. While we restrict acceleration measurements to KompeteAI due to tight pipeline integration, the proposed paradigms are general and may be adapted to other architectures with minimal conceptual changes.
Figure 2: Comparison of our pipeline with AIDE, RD-agent and ML-Master on Contemporary and MLE-Bench parts of Kompete-bench. All systems use gemini-2.5-flash as the underlying LLM, except for ML-Master, which uses deepseek-r1. Each was run 3 times with different seeds, results are averaged, and each run was limited to 6 hours.
@@ 6 Conclusions
We introduced KompeteAI, a new multi-agent AutoML framework, and demonstrated its strong performance on diverse and challenging tasks. Our contributions include a state-of-the-art architecture with dynamic integration of external knowledge sources and novel tree-based operators for adding and merging, combining LLM decision-making with algorithmic methods. We further propose an Accelerated Evaluation and Debugging Paradigm to address execution bottlenecks, applicable to general LLM-based AutoML architectures. Finally, we show that the current benchmark suffers from significant bias and introduce Kompete-Bench to overcome this limitation.
Acknowledgments.
The research has been supported by the Ministry of Economic Development of the Russian Federation (agreement No. 139-10-2025-034 dd. 19.06.2025, IGK  [^00] C313925P4D0002).
Appendix A Future work
Our work opens several promising directions for future research.
@@ Appendix A Future work
Appendix B Algorithms
B.1 Adding
@@ B.2 Merging
Appendix C Kompete-bench
0-99 Teams 100-249 Teams 250-999 Teams 1000+ Teams
Bronze Top 40% Top 40% Top 100 Top 10%
Silver Top 20% Top 20% Top 50 Top 5%
Gold Top 10% Top 10 Top 10 + 0.2%* Top 10 + 0.2%*
Table 5: Thresholds for winning a medal in Kaggle competitions vary depending on the number of teams participating in each competition. We implement the same thresholds as in MLE-bench. *the threshold increases by 1 for every 500 additional teams. Source: Kaggle (2024).
@@ 0-99 Teams 100-249 Teams 250-999 Teams 1000+ Teams
Figure A3: End dates for competitions presented in Kompete-Bench
@@
Appendix D Setup details
@@
KompeteAI. The list of hyperparameters used when running KompeteAI on benchmarks is given in Table 6.The time_run_minutes parameter sets the maximum runtime for the entire multi-agent system in minutes, after which the process will terminate. The runtime_error_time defines the time limit (in minutes) after which a generated code will be stopped. The subset_size_in_percent specifies the percentage of the dataset to be used for quick validation. The validator_size_threshold sets a threshold for the dataset size; if the data exceeds this value, a subset is used for training and validation. The number_of_ideas_eda determines how many exploratory data analysis (EDA) ideas are generated per iteration. Similarly, number_of_ideas_data and number_of_ideas_modelling control the number of ideas related to feature engineering and model training. The max_add_idea parameter limits how many new ideas can be added to the idea pool in a single adding iteration. The number_of_selected_node specifies how many nodes are selected for adding expansion at each step. The number_of_iterations_parents sets how many iterations parent nodes participate in generating new ideas, while number_of_selected_node_merging determines how many nodes are chosen for merging at each iteration. The number_of_ideas_min and number_of_ideas_max define the minimum and maximum number of ideas, which are used as anchor ideas for the scoring model. The retrieve_n_papers and retrieve_n_competitions parameters control how many papers are retrieved from arXiv papers and Kaggle solutions. The number_rag_ideas sets how many ideas are generated using RAG. The memory_size parameter determines how many recent ideas, solutions, or states each agent remembers for learning and decision-making. Alternatively, if memory_size is set to nearest_nodes, the agent’s memory consists of the most similar nodes or ideas, rather than a fixed number, allowing for more contextually relevant recall.
@@
RD-Agent. The list of hyperparameters used when running RD-agent on benchmarks is given in Table 7. The debug_timeout parameter sets the maximum time, in seconds, that is allowed to debug one generated code. The full_timeout parameter defines the overall time limit for the system. The if_action_choosing_based_on_UCB flag determines whether the agents select their actions using the Upper Confidence Bound (UCB) strategy. The enable_knowledge_base flag indicates whether a shared knowledge base is enabled for the agents. The loop_n parameter specifies the number of main iterations the system will execute, set to 2000 to avoid early stopping not by the time limit. Finally, the max_trace_num parameter limits the number of traces that can be created during system execution.
@@
AIDE.
The list of hyperparameters used when running RD-agent on benchmarks is given in Table 8. The steps parameter defines the maximum number of steps the entire system can perform during a run. The max_debug_depth specifies the maximum depth for recursive code debugging. The debug_prob parameter suggests that debugging is enabled for every generated code, meaning that all relevant information will be recorded without any sampling. Finally, the time_limit parameter represents the maximum allowed wall-clock time (in seconds) for the experiment.
@@
Appendix E Metrics
Competition RD-agent AIDE ML-Master MLE-STAR KompeteAI aerial-cactus-identification 58 5 58 88 79 denoising-dirty-documents 3 None 15 0 12 dog-breed-identification 37 62 54 44 58 dogs-vs-cats-redux-kernels-edition 86 6 47 84 91 jigsaw-toxic-comment-classification-challenge 40 26 47 30 88 leaf-classification 33 24 35 54 46 mlsp-2013-birds None 0 9 0 None nomad2018-predict-transparent-conductors 78 33 76 30 22 plant-pathology-2020-fgvc7 12 62 45 38 68 random-acts-of-pizza 58 47 56 58 70 spooky-author-identification 67 47 60 58 63 tabular-playground-series-dec-2021 17 22 26 13 29 tabular-playground-series-may-2022 16 32 29 23 43 text-normalization-challenge-english-language 3 9 25 7 27 text-normalization-challenge-russian-language 17 14 36 33 18